OpenStudy (anonymous):

how would I expand log7 49xyz

OpenStudy (yttrium):

same as: log7 49 + log7 x + log7 y + log7 z = 2 + log7 x + log7 y + log7 z

OpenStudy (anonymous):

Where does the 2 come from I know it has to do something with 7^2=49 but I dont understand how to get it

OpenStudy (yttrium):

do you know that log(a) b = c is same as a^c = b In this case, we'll have log7 49 = 2 because 7^2 = 49

OpenStudy (anonymous):

Oh i get it thanks